# Media Bias Detector: Designing and Implementing a Tool for Real-Time Selection and Framing Bias Analysis in News Coverage

### v1

@@ 5.3. Quantitative Measures
We used a mixed-methods approach to evaluate the Media Bias Detector. In addition to collecting qualitative feedback on how participants used our tool, we also measured key dimensions of the tool’s impact: bias identification and awareness (RQ1), comparative analysis (RQ1), and user engagement (RQ3). To quantify these aspects, we asked participants to rate their agreement on a 7-point Likert-scale (1 = Strongly Disagree, 7 = Strongly Agree), similar to prior work (Zavolokina et al., 2024; Reza et al., 2024). Furthermore, to address RQ1, we assessed subjective workload using the NASA-TLX procedure with weighting (Hart, 2006), a method commonly used in HCI research to measure a user’s cognitive demand when performing a task.
@@ 5.4. Procedure
During the pre-task interview, we explored participants’ general media consumption habits, preferred news sources, and their engagement with news and politics. We also asked for their views on how news outlets may emphasize certain topics while overlooking others (selection bias) and how the tone and language of articles can influence readers’ perceptions (framing bias). Next, participants undertook the sequential 10-minute media bias comparative analysis tasks. Participants were encouraged to think aloud as they performed both tasks (Ramey et al., 2006; Nielsen et al., 2002). Before the Media Bias Detector task, we showed a brief walkthrough video highlighting its main features. At the end of each task, we provided participants with a survey link to complete the evaluation questions, including Likert-scale and NASA-TLX questions.
@@ 5.5. Analysis
We collected interview transcripts, observation notes, and a set of quantitative responses for each task. We coded and analyzed the qualitative data using Miro (Miro, 2024), which allowed us to collaboratively code the data using digital sticky notes (Johnson, 2022; Burgess et al., 2021; Geraci et al., 2014). Four team members manually reviewed the interview transcripts noting key points. From these notes, we used an inductive coding approach to generate initial descriptive codes and identify relationships between themes discussed in participant quotes (Braun and Clarke, 2006). Subsequently, our team reviewed the codes and preliminary themes with the assistance of an AI summarization tool and had multiple group discussions to note similarities and differences before agreeing on the main broad themes we learned from our expert interviews (McDonald et al., 2019).
@@ 6.3. RQ2: Challenges Surrounding Trust and Transparency in LLM-Driven Bias Detection
Several participants suggested that adding features to increase transparency could help build trust (P1, P2, P9). P1 expressed, ”For me, someone who is a bit more skeptical of AI research in general, being able to see the articles [labeled under a topic] would make me feel better”. P2 also suggested that having access to article texts would help them evaluate how the LLM processes articles with mixed tones, particularly in climate change coverage, where optimistic, positive outlooks may exist within negative stories (Ojala et al., 2021). This feature would allow them to determine if the LLM captures subtle shifts in tone or oversimplifies it as neutral.
@@ 8. Discussion
From a user experience perspective, we chose to distill complex information into two salient outcome measures that vary significantly by article within publisher: lean and tone, where we emphasize that tone is unrepresented in existing bias tools but is increasingly considered important to audience perception (Robertson et al., 2023). The Media Bias Detector focuses on these two dimensions to provide an accessible, user-friendly starting point before expanding to additional variables. Several participants noted that the tool’s interactivity gave users more autonomy to form their own judgments. Participants appreciated the dual focus on lean and tone, as well as the affordances and customization options offered by the tool. However, some participants suggested that simplifying the tool could improve usability, which implied that widening the scope of the Media Bias Detector to additional dimensions could introduce too much complexity in learning how to use it. These findings demonstrate that while deconstructing the multidimensional aspects of media bias is both feasible and impactful, user-centered design is crucial. The Media Bias Detector contributes a design-centered approach to media bias tools, focusing on interactive evaluation across multiple perspectives, within categories, and while comparing sources to each other.
@@
Our approach includes humans in the loop to validate the model’s outputs and ensure they align with human judgment. This hybrid system balances automation with accountability, maintaining consistency without placing sole reliance on automated processes. Encouragingly, both experts and everyday news consumers generally showed increased trust in the tool after learning about the human-in-the-loop process. This finding is consistent with prior research showing that trust in AI grows when people know it has higher accuracy (Zhang et al., 2020). In our case, the humans-in-the-loop act as accuracy validators by ensuring the model’s outputs are reasonable and reliable.
@@
Beyond immediate trust, our findings highlight broader issues associated with human-AI tool development. While experts who work closely with language models and understand their faulty nature remain skeptical, regular users are quickly adjusting to the new normal with LLMs. Prior research has shown that trust in AI grows with familiarity (Yu et al., 2017). Assuming the research community continues to validate these models and ensure they remain calibrated, the trust growing among users is encouraging. However, it also underscores the critical need to design tools that keep humans in the loop, prioritize transparency, and incorporate safeguards to address initial skepticism—preventing potential failures before it’s too late.
@@
We also discovered that news consumers saw the Media Bias Detector as a tool that could shift their views on media bias over time, with many expressing interest in additional educational resources. This suggests that the tool not only revealed biases they hadn’t noticed before but also sparked curiosity to learn more. In our follow-up study, a significant number of participants indicated they would benefit from extra guidance on using the tool effectively. This suggests that incorporating more resources on communication techniques like prebunking (Amazeen et al., 2022), accuracy nudges and fact-checking (Pennycook and Rand, 2022), and debunking (Chan et al., 2017) could supplement the Media Bias Detector and further support the learning process. Although accuracy nudges and fact-checking are more pertinent to social media content and may not be directly applicable to mainstream media outlets—which typically do not contain outright factual inaccuracies—parallel strategies could be developed to prompt news consumers to recognize subtle biases, such as those involving framing and selection. The Media Bias Detector is purposefully designed to be modular, prioritizing popular news publishers with the greatest reach (Allen et al., 2024). By focusing on high-impact publishers, we can sustainably validate LLMs’ outputs and implement necessary safeguards before expanding the tool further.
@@ 9.2. Data-Related Limitations
The Media Bias Detector currently tracks only ten publishers and collects their top 20 articles in every interval. While this covers the most prominent news published by the most widely-read publishers, it excludes coverage from local or more niche sources. Several participants highlighted that this approach misses other perspectives in the news. We built the dashboard with ten publishers, due to the significant financial and engineering costs associated with data collection and labeling pipelines. Each publisher requires a tailored data collection pipeline which involves manual text preprocessing. Additionally, using state-of-the-art, closed-source models (i.e., OpenAI’s GPT-4o) on a daily basis incurs significant costs that scale linearly with the number of publishers analyzed. To select these ten publishers, we sought out a set of sources that are popular among news readers (Statista, 2023), while also maintaining diversity across the political spectrum (we added Breitbart for its agenda-setting power on the right, The Guardian and Huffpost on the left, and the Wall Street Journal for its financial leadership). We are currently in the process of expanding our data collection pipeline to cover the top 30 stories per news publisher and plan to expand our list of publishers beyond legacy media.
Another limitation of our tool is that the topic and subtopic lists are driven by what is actively discussed in the news. When the human-in-the-loop annotators identify major themes that have emerged in the news, we make updates to the topic list. This creates a blind spot for important issues that are not being reported by research assistants, as they may never appear in our analysis. Additionally, our current approach assigns each article to a single topic and subtopic. This rigid assignment of topics and subtopics can overlook articles that fall under multiple options. Not including them in one of those options affects what we see in the coverage numbers and may not fully reflect reality. This highlights the broader challenge of classifying a complex news article into a finite number of categories. In future iterations of the Media Bias Detector, we may explore more flexible methods to accommodate overlapping themes. For example, allowing a single article to be associated with ”most likely” topics or subtopics or adjusting prompts to support multiclass classification.
@@ 9.3. Limitations of LLMs in Media and Politics
Our tool relies on instruction-tuned, pre-trained LLMs to generalize tasks in a zero-shot setting, which introduces inherent challenges (Zhang et al., 2023b). While LLMs are efficient for text classification, they are not immune to biases present in their training data, which can subtly influence downstream tasks such as political lean labeling (Feng et al., 2023). This has important implications for the Media Bias Detector’s results, particularly when analyzing political news topics. When prompting the LLM for explanations on its political lean label, we can identify that they draw conclusions from not only the article text itself, but from context and inferences from its training data. This results in the LLM inherently associating certain topics with Democratic leanings (such as climate change or women’s healthcare), and others with Republican leanings (such as immigration or crime), regardless of the specific arguments presented in the article. We must be cautious as the nature of these contexts and inferences may shift over time.
When measured using both novel and established methods to assess political bias, GPT model variants, including GPT-4 and GPT-3.5-Turbo, tend to align with a left-leaning political ideology (Feng et al., 2023; Rutinowski et al., 2023; Motoki et al., 2024). This may relate to the model’s tendency to align certain issues with the Democratic or Republican party’s political agenda. While individual political lean labels could be subject to LLM biases, their relative values, and the differences between labels on articles about the same event, are informative. Our research is grounded in comparative analysis between publishers on multiple levels, and we intentionally provide granularity in all our analyses to allow users to draw their own conclusions. Furthermore, the Media Bias Detector emphasizes features like tone, news topics, and event coverage, which are less susceptible to political bias in LLMs compared to political lean labels.
@@ 10. Conclusion
We believe that helping people identify and quantify bias across different stories and publishers, along with exposing readers to alternative narratives, can enhance public awareness of the editorial choices made by news publishers. From this perspective, the Media Bias Detector could help counter the potential for creating separate realities by empowering users to independently compare how different publishers cover the same topics or events (Bram, 2024). For instance, participants in our user study noted that the Events dashboard helped them see discrepancies in coverage of major events, such as the Gaza ceasefire negotiations, where some outlets prioritized domestic political stories instead. Additionally, everyday users reported that the ability to drill down into subtopics and compare political lean and tone encouraged them to challenge their preconceptions about specific publishers. By providing a high-level view of the media landscape, the Media Bias Detector can help users engage with diverse perspectives and hold media outlets accountable for their power to shape public perception. Moreover, the tool may also serve as a valuable resource for researchers, educators, and journalists, enabling them to check their own biases and quantitatively assess media coverage.
We close by re-emphasizing that the goal of the Media Bias Detector is to empower users to discover bias on their own, not to tell them what is biased and what is not. Our focus is on understanding how news publishers’ decisions influence public perception, and how emerging narratives are shaped by the emphasis placed on them by different mainstream media outlets. This exploration is not about finding fault, but about analyzing how the natural processes of news selection and framing influence the stories we encounter in our everyday lives (Watts et al., 2021). We therefore hope our work contributes to meaningful discussions at the intersection of media, the public, and politics, by promoting critical thinking about how news is produced, consumed, and absorbed, and ultimately the worldviews and choices it shapes.
@@ Acknowledgements.
We thank Polygraph for their support in developing the frontend. We also thank Ajay Patel and Bryan Li for helpful discussions on the methodology. We also express our gratitude to Harsh Kumar for encouraging us to conduct user studies on the Media Bias Detector and present it to an HCI audience. This research was developed with funding from Richard Jay Mack and the Defense Advanced Research Projects Agency’s (DARPA) SciFy program (Agreement No. HR00112520300). The views expressed are those of the authors and do not reflect the official policy or position of the Department of Defense or the U.S. Government.

### v2

@@ 5.3. Quantitative Measures
We used a mixed-methods approach to evaluate the Media Bias Detector. In addition to collecting qualitative feedback on how participants used our tool, we also measured key dimensions of the tool’s impact: bias identification and awareness (RQ1), comparative analysis (RQ1), and user engagement (RQ3). To quantify these aspects, we asked participants to rate their agreement on a 7-point Likert-scale (1 = Strongly Disagree, 7 = Strongly Agree), similar to prior work (think_fast; reza_abscribe). Furthermore, to address RQ1, we assessed subjective workload using the NASA-TLX procedure with weighting (hart2006nasa), a method commonly used in HCI research to measure a user’s cognitive demand when performing a task.
@@ 5.4. Procedure
During the pre-task interview, we explored participants’ general media consumption habits, preferred news sources, and their engagement with news and politics. We also asked for their views on how news outlets may emphasize certain topics while overlooking others (selection bias) and how the tone and language of articles can influence readers’ perceptions (framing bias). Next, participants undertook the sequential 10-minute media bias comparative analysis tasks. Participants were encouraged to think aloud as they performed both tasks (ramey_2006; nielsen2002getting). Before the Media Bias Detector task, we showed a brief walkthrough video highlighting its main features. At the end of each task, we provided participants with a survey link to complete the evaluation questions, including Likert-scale and NASA-TLX questions.
@@ 5.5. Analysis
We collected interview transcripts, observation notes, and a set of quantitative responses for each task. We coded and analyzed the qualitative data using Miro (miro), which allowed us to collaboratively code the data using digital sticky notes (johnson_miro_2022; burgess2021sticky; geraci2014work). Four team members manually reviewed the interview transcripts noting key points. From these notes, we used an inductive coding approach to generate initial descriptive codes and identify relationships between themes discussed in participant quotes (braun2006using). Subsequently, our team reviewed the codes and preliminary themes with the assistance of an AI summarization tool and had multiple group discussions to note similarities and differences before agreeing on the main broad themes we learned from our expert interviews (mcdonald_2019).
@@ 6.3. RQ2: Challenges Surrounding Trust and Transparency in LLM-Driven Bias Detection
Several participants suggested that adding features to increase transparency could help build trust (P1, P2, P9). P1 expressed, ”For me, someone who is a bit more skeptical of AI research in general, being able to see the articles [labeled under a topic] would make me feel better”. P2 also suggested that having access to article texts would help them evaluate how the LLM processes articles with mixed tones, particularly in climate change coverage, where optimistic, positive outlooks may exist within negative stories (ojala2021anxiety). This feature would allow them to determine if the LLM captures subtle shifts in tone or oversimplifies it as neutral.
@@ 8. Discussion
From a user experience perspective, we chose to distill complex information into two salient outcome measures that vary significantly by article within publisher: lean and tone, where we emphasize that tone is unrepresented in existing bias tools but is increasingly considered important to audience perception (robertson2023negativity). The Media Bias Detector focuses on these two dimensions to provide an accessible, user-friendly starting point before expanding to additional variables. Several participants noted that the tool’s interactivity gave users more autonomy to form their own judgments. Participants appreciated the dual focus on lean and tone, as well as the affordances and customization options offered by the tool. However, some participants suggested that simplifying the tool could improve usability, which implied that widening the scope of the Media Bias Detector to additional dimensions could introduce too much complexity in learning how to use it. These findings demonstrate that while deconstructing the multidimensional aspects of media bias is both feasible and impactful, user-centered design is crucial. The Media Bias Detector contributes a design-centered approach to media bias tools, focusing on interactive evaluation across multiple perspectives, within categories, and while comparing sources to each other.
@@
Our approach includes humans in the loop to validate the model’s outputs and ensure they align with human judgment. This hybrid system balances automation with accountability, maintaining consistency without placing sole reliance on automated processes. Encouragingly, both experts and everyday news consumers generally showed increased trust in the tool after learning about the human-in-the-loop process. This finding is consistent with prior research showing that trust in AI grows when people know it has higher accuracy (zhang_accuracy). In our case, the humans-in-the-loop act as accuracy validators by ensuring the model’s outputs are reasonable and reliable.
@@
Beyond immediate trust, our findings highlight broader issues associated with human-AI tool development. While experts who work closely with language models and understand their faulty nature remain skeptical, regular users are quickly adjusting to the new normal with LLMs. Prior research has shown that trust in AI grows with familiarity (yu_user_trust). Assuming the research community continues to validate these models and ensure they remain calibrated, the trust growing among users is encouraging. However, it also underscores the critical need to design tools that keep humans in the loop, prioritize transparency, and incorporate safeguards to address initial skepticism—preventing potential failures before it’s too late.
@@
We also discovered that news consumers saw the Media Bias Detector as a tool that could shift their views on media bias over time, with many expressing interest in additional educational resources. This suggests that the tool not only revealed biases they hadn’t noticed before but also sparked curiosity to learn more. In our follow-up study, a significant number of participants indicated they would benefit from extra guidance on using the tool effectively. This suggests that incorporating more resources on communication techniques like prebunking (amazeen2022cutting), accuracy nudges and fact-checking (pennycook2022nudging), and debunking (chan2017debunking) could supplement the Media Bias Detector and further support the learning process. Although accuracy nudges and fact-checking are more pertinent to social media content and may not be directly applicable to mainstream media outlets—which typically do not contain outright factual inaccuracies—parallel strategies could be developed to prompt news consumers to recognize subtle biases, such as those involving framing and selection. The Media Bias Detector is purposefully designed to be modular, prioritizing popular news publishers with the greatest reach (allen2024quantifying). By focusing on high-impact publishers, we can sustainably validate LLMs’ outputs and implement necessary safeguards before expanding the tool further.
@@ 9.2. Data-Related Limitations
[…]
@@ 9.3. Limitations of LLMs in Media and Politics
Our tool relies on instruction-tuned, pre-trained LLMs to generalize tasks in a zero-shot setting, which introduces inherent challenges (zhang2023instruction). While LLMs are efficient for text classification, they are not immune to biases present in their training data, which can subtly influence downstream tasks such as political lean labeling (feng-etal-2023-pretraining). This has important implications for the Media Bias Detector’s results, particularly when analyzing political news topics. When prompting the LLM for explanations on its political lean label, we can identify that they draw conclusions from not only the article text itself, but from context and inferences from its training data. This results in the LLM inherently associating certain topics with Democratic leanings (such as climate change or women’s healthcare), and others with Republican leanings (such as immigration or crime), regardless of the specific arguments presented in the article. We must be cautious as the nature of these contexts and inferences may shift over time.
When measured using both novel and established methods to assess political bias, GPT model variants, including GPT-4 and GPT-3.5-Turbo, tend to align with a left-leaning political ideology (feng-etal-2023-pretraining; rutinowski2023self; motoki2024more). This may relate to the model’s tendency to align certain issues with the Democratic or Republican party’s political agenda. While individual political lean labels could be subject to LLM biases, their relative values, and the differences between labels on articles about the same event, are informative. Our research is grounded in comparative analysis between publishers on multiple levels, and we intentionally provide granularity in all our analyses to allow users to draw their own conclusions. Furthermore, the Media Bias Detector emphasizes features like tone, news topics, and event coverage, which are less susceptible to political bias in LLMs compared to political lean labels.
@@ 10. Conclusion
[…]
We close by re-emphasizing that the goal of the Media Bias Detector is to empower users to discover bias on their own, not to tell them what is biased and what is not. Our focus is on understanding how news publishers’ decisions influence public perception, and how emerging narratives are shaped by the emphasis placed on them by different mainstream media outlets. This exploration is not about finding fault, but about analyzing how the natural processes of news selection and framing influence the stories we encounter in our everyday lives (Watts_etal). We therefore hope our work contributes to meaningful discussions at the intersection of media, the public, and politics, by promoting critical thinking about how news is produced, consumed, and absorbed, and ultimately the worldviews and choices it shapes.
@@ Acknowledgements.
[…]